OpenStudy (braves12):

A cone with a radius of 3" has a total area of 24 sq. in. Find the volume of the cone. 24 15 12

OpenStudy (anonymous):

24pi=pi*9+pi*3*l l=5 l^2=r^2+h^2 5^2=3^2+h^2 h=4

OpenStudy (anonymous):

Volum=1/3*pi*r^2*h V=1/3*3.4*3^2*4 V=37.7

OpenStudy (anonymous):

V=1/3 pi * r^2 * h=80